OpenStudy (anonymous):

So much gratitude for help w this: A bag contains 2 steel balls and 5 brass balls. The total weight is 13 pounds. If 2 steel balls are added and 2 brass balls are removed, the weight decreases to 12 pounds. How much does each kind of ball weigh? A. steel, 4 pounds; brass, 1 poundB. steel, 1.5 pounds; brass, 2 poundsC. steel, 0.75 pounds; brass, 3 poundsD. steel, 3 pounds; brass, 4 pounds.

OpenStudy (kagıtucak):

let a stell ball be x pounds and a brass ball be y pounds. From the first and second sentence: 2x+5y=13 From the third sentence: 4x+3y=12 Multiply first equation with -2 and sum with second equation -10y+3y=-14 -7y=-14 y=2 Plug this into the first equation 2x+10=13 x=1.5 So it is B stell ball is 1.5 pounds and brass ball is 2 pounds.